A gun of mass 'M' fires a bullet of mass 'm' with a velocity 'v' relative to the gun. The average force required to bring the gun to rest in 0.5 seconds is

Grade:11

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ramya

conservation of linear momentum

let the velocity of gun is V1 wrt earth

so velocity of bullet wrt eart is v-V1

m(v-v1) =Mv1

  v1=mv/(m+M)

now distance travel by gun in 0.5 sec

0=v1-a*0.5

a= 2v1     wher a is retardation of gun

so S= v1 *0.5 -1/2 *a *(0.5)2

       =v1 /4

now from work energy theorem total work done is equal to change in kinetic energy

F.S =1/2 M *v12    -0

F.v1 /4 = 1/2 M *v12

F=2Mv1

  =2Mmv/(m+M)


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

Ishana
6 years ago
This is not the correct answer for this question. I asked you to find total kinetic energy in the question but you have found something else.????????!!!!!!!!!!!